# Multiple dispersive bounds. I) An improved z-expansion

### v2

@@ III Multiple dispersive bounds
The quantity $\chi[f]$ is by definition an inclusive (positive) quantity, since all intermediate on-shell states $\overline{H}_{1}H_{2}$ above the pair-production threshold $t_{+}=(m_{1}+m_{2})^{2}$ are summed up. The second novel ingredient, proposed in this work, is the introduction of a set of limited, real and non-negative kernels $\widetilde{K}_{p}(z)$ (where p=1,,Pp=1,...,P), constrained by
@@ III.2 Evaluation of the multiple dispersive bounds $\chi_{p}^{U}$
We now briefly describe the last option, in which the kinematical function $\phi(z)$ is the outer function given by Eq. (12). To this end, we introduce a set of limited, real and non-negative kernels $K_{p}(\tau)$ with p=1,2,Pp=1,2,...P given in terms of the Euclidean time-distance $\tau$. Again, the choice of such kernels is quite generic. In addition to the case of the step functions, an interesting choice is represented by smoothed time windows. A recent example is provided by the short-, intermediate- and long-distance windows introduced by the RBC Collaboration [43] for investigating on the lattice the Hadronic Vacuum Polarization(HVP) contribution to the muon g2g-2. Recalling Eq. (16), we use the kernel $K_{p}(\tau)$ to define the following dispersive bounds
@@ IV Conclusions
In this work we have proposed the introduction of two novel ingredients in the implementation of the widely used BGL $z$-expansion of hadronic form factors.
@@
On the other hand side, we have proposed the use of multiple dispersive bounds instead of a single, total dispersive bound through the introduction of a set of appropriate kernel functions in the evaluation of the hadronic susceptibilities. In this work we have focused on the general implementation of multiple dispersive bounds in the BGL framework in the absence of sub-threshold branch-cuts. The numerical applications of multiple dispersive bounds depend strongly on the specific physical process of interest and, therefore, they will be left to future, dedicated works. The generalization of multiple dispersive bounds to the case in which sub-threshold branch-cuts are present, is investigated in the companion paper [6], where an extensive numerical application of double dispersive bounds is presented.

### v3

@@ III Multiple dispersive bounds
The quantity $\chi[f]$ is by definition an inclusive (positive) quantity, since all intermediate on-shell states $\overline{H}_{1}H_{2}$ above the pair-production threshold $t_{+}=(m_{1}+m_{2})^{2}$ are summed up. The second ingredient, proposed in this work, is the introduction of a set of limited, real and non-negative kernels $\widetilde{K}_{p}(z)$ (where p=1,,Pp=1,...,P), constrained by
@@ III.2 Evaluation of the multiple dispersive bounds $\chi_{p}^{U}$
We now briefly describe the last option, in which the kinematical function $\phi(z)$ is the outer function given by Eq. (13). To this end, we introduce a set of limited, real and non-negative kernels $K_{p}(\tau)$ with p=1,2,Pp=1,2,...P given in terms of the Euclidean time-distance $\tau$. Again, the choice of such kernels is quite generic. In addition to the case of the step functions, an interesting choice is represented by smoothed time windows. A recent example is provided by the short-, intermediate- and long-distance windows introduced by the RBC Collaboration [46] for their studies of the Hadronic Vacuum Polarization (HVP) contribution to the muon g2g-2. Recalling Eq. (18), we use the kernel $K_{p}(\tau)$ to define the following dispersive bounds
@@ IV Conclusions
In this work we have proposed the introduction of two ingredients in the implementation of the widely used BGL $z$-expansion of hadronic form factors.
@@
On the other hand side, we have proposed the use of multiple dispersive bounds instead of a single, total dispersive bound through the introduction of a set of appropriate kernel functions in the evaluation of the hadronic susceptibilities. In this work we have focused on the implementation of multiple dispersive bounds in the BGL framework in the absence of sub-threshold branch-cuts. The numerical applications of multiple dispersive bounds depend strongly on the specific physical process of interest and, therefore, they will be left to future, dedicated works. The application of our proposal for exploiting multiple dispersive bounds to the case in which sub-threshold branch-cuts are present, is investigated in the companion paper [6], where an extensive numerical application of double dispersive bounds is presented.